A certain substance X condenses at a temperature of 147.2 °C. But if a 800. g sample of X is prepared with 77.9 g of urea ((NH2)2Co) dissolved in it, the

sample is found to have a condensation point of 150.7 °C instead. Calculate the molal boiling point elevation constant Kb of X.

Be sure your answer has the correct number of significant digits 



1
Expert's answer
2021-03-04T06:43:25-0500

Mass of urea= 77.9g

Molar mass of urea= 60g/mol

Moles of urea= mass/molar mass = 77.9/60= 1.298mol

Molality of urea=moles of urea/Kg of solvent

= 1.298mol/0.800Kg

= 1.623Molal


Now, we can now find the boiling point elevation constant Kb by using the relation:

∆Tb= Kbm

Where ∆Tb is the boiling point elevation, m is the molality of solute.

∆Tb= T2-T1= Kbm

150.7°C - 147.2°C = Kb x 1.623

Kb= 3.5/1.623=2.157